# Towards Robust Influence Functions with Flat Validation Minima

### v2

@@ B.3 Text-To-Image Generation
For style generation, we fine-tune the model by appending a style description to the text prompt, instructing it to generate images in a specific style. We use 200 training image-text pairs and 50 validation image-text pairs, resulting in a total of 600 training data points and 150 validation data points. We use the DreamBooth dataset (Ruiz et al., 2023), which contains 30 different subjects The text follows the format: “Generate an image in a specific custom style. text-data”, where custom is replaced with either “cartoon”, “pixelart”, or “black and white line sketch”, and text-data is replaced with a text sequence from the training dataset. A detailed description of the style prompts is provided in Table 9.
For subject generation, we use Google’s DreamBooth dataset (Ruiz et al., 2023), which contains 31 unique subjects across categories such as backpack, dog, bowl, and sneaker. Each subject has four to six images, with three examples per subject used for training and the remaining for validation. For each subject, 3 data points are used for the training dataset and 1 to 3 data points are used for the validation dataset. To enable the model to distinguish between different subjects, we assign a unique random string to each subject in the prompt. For instance, two different dogs are referenced as “a i0VpE dog” and “a oHwLM dog” in the training and validation sets, respectively.
Training. We use LoRA (Hu et al., 2022) to fine-tune stable-diffusion-v1.5 (Rombach et al., 2022). We apply LoRA to every attention layer in the stable-diffusion-v1.5 model. The basic hyper-parameters settings are listed as follows: minibatch size (4), optimizer (AdamW (Loshchilov & Hutter, 2019)), initial learning rate ($10^{-4}$), number of steps (10000), and learning rate decay (cosine). For LoRA hyperparameters, we set $r=2$ and $\alpha=2$. The training was performed using the HuggingFace Peft library (Mangrulkar et al., 2022).
@@ C.1 Comparison with EK-FAC
We compare our proposed VM/FVM approach with EK-FAC (Grosse et al., 2023) on mislabeled sample detection tasks, following the experimental setup described in Section B.1. The results are presented in Table 10, where * indicates performance based on the best training checkpoints. As shown, our method consistently outperforms the EK-FAC-based influence function in identifying mislabeled data.
@@ C.3 Sharpness-Aware Optimizer
We compare three different sharpness-aware optimizer, including SAM (utilized in our initial approach), ASAM (Kwon et al., 2021), and F-SAM (Li et al., 2024), on CIFAR-10N Worst and CIFAR-100N Noisy datasets. The results (ROC AUC) are presented in Table 12. As observed, different sharpness-aware optimizers do affect the final results, with F-SAM achieving the best performance. This supports our hypothesis that better flat validation minima can lead to more accurate influence function estimation.
@@ C.4 Inverse Hessian Approximation
To quantitatively discover the impact of the approximation on the inverse Hessian, we conduct experiments on the mislabel detection task. Specifically, we replace the computation of the inverse Hessian $\tilde{H}_{val}^{-1}$ in Equation 21 from diagonal Fisher to LiSSA (Koh & Liang, 2017) and DataInf (Kwon et al., 2024).
@@ E.1 The Parameter Change $\tilde{\theta}_{z_{\text{tr}}}-\tilde{\theta}$
Here, for completeness, we derive the parameter change $\tilde{\theta}_{z_{\text{tr}}}-\tilde{\theta}$ in the context of loss minimization. This derivation builds upon the analysis of $\theta^{\star}_{z_{\text{tr}}}-\theta^{\star}$ presented in Koh & Liang (2017).
@@ F.1 Influence Function
The use of influence-based diagnostics dates back to seminal works such as Cook & Weisberg (1982). More recently, Koh & Liang (2017) introduced Influence Functions to large-scale Deep Neural Networks (DNNs), sparking significant research interest. However, subsequent studies have identified challenges in applying IF to DNNs. Basu et al. (2021) highlighted the fragility of IF in deep learning scenarios, while Bae et al. (2022) demonstrated that practical influence function estimates often fail to align with leave-one-out retraining in nonlinear networks. Instead, they approximate a different quantity known as the proximal Bregman response function (PBRF).
Still, IF has inspired a wave of research aimed at improving its computational efficiency and scalability for large models. TracIn (Pruthi et al., 2020) improves computational efficiency by tracking loss changes using gradient information from multiple training checkpoints. FastIF (Guo et al., 2021) accelerates influence estimation by selecting the $k$-nearest training samples around a given validation sample. Schioppa et al. (2022) speed up the inverse Hessian calculation based on Arnoldi iteration (ARNOLDI, 1951). Grosse et al. (2023) proposed Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) to scale Influence Function computations to large language models (LLMs). GEX (Kim et al., 2023) alleviates the bilinear constraint of standard Influence Functions and leverages Geometric Ensemble (GE) for more precise influence estimation. DataInf (Kwon et al., 2024) further enhanced efficiency by deriving a closed-form influence estimation, making it particularly effective for parameter-efficient fine-tuning techniques such as LoRA (Hu et al., 2022).
@@ F.2 Sharpness-Aware Minimization
The relationship between the flatness of local minima and generalization has been extensively explored in prior works (Dinh et al., 2017; Li et al., 2018). More recently, numerous studies have aimed to improve model generalization by explicitly seeking flatter minima (Chaudhari et al., 2017; Tsuzuku et al., 2020; Foret et al., 2021). Among these, Sharpness-Aware Minimization (SAM) (Foret et al., 2021) has attracted considerable attention. SAM introduces a general training framework that formulates optimization as a min-max problem, encouraging model parameters to lie in neighborhoods with uniformly low loss, thereby achieving state-of-the-art generalization across a wide range of tasks. Building on SAM, subsequent works have advanced the approach by refining the geometric definition of the neighborhood (Kwon et al., 2021; Kim et al., 2022), proposing improved surrogate loss functions (Zhuang et al., 2022), developing more adaptive adversarial strategies (Li & Giannakis, 2023), and improving training efficiency (Jiang et al., 2023).
@@ G.1 Remark on the Sign of Influence
In previous studies, the same training sample can yield positive or negative influence depending on the paper, due to differing definitions of loss change. This ambiguity arises particularly between $\ell(z,\theta_{z})-\ell(z,\theta)$ and $\ell(z,\theta)-\ell(z,\theta_{z})$. For example, Koh & Liang (2017) define the influence as $\ell(z,\theta_{z})-\ell(z,\theta)$, whereas Pruthi et al. (2020) define it as $\ell(z,\theta)-\ell(z,\theta_{z})$. In this paper, we adopt the previous approach, measuring loss change as $\ell(z,\theta_{z})-\ell(z,\theta)$.

### v3

@@ B.3 Text-To-Image Generation
For style generation, we fine-tune the model by appending a style description to the text prompt, instructing it to generate images in a specific style. We use 200 training image-text pairs and 50 validation image-text pairs, resulting in a total of 600 training data points and 150 validation data points. We use the DreamBooth dataset (DBLP:conf/cvpr/RuizLJPRA23:DreamBooth), which contains 30 different subjects The text follows the format: “Generate an image in a specific custom style. text-data”, where custom is replaced with either “cartoon”, “pixelart”, or “black and white line sketch”, and text-data is replaced with a text sequence from the training dataset. A detailed description of the style prompts can be found in DBLP:conf/iclr/KwonWW024:DataInf.
For subject generation, we use Google’s DreamBooth dataset (DBLP:conf/cvpr/RuizLJPRA23:DreamBooth), which contains 31 unique subjects across categories such as backpack, dog, bowl, and sneaker. Each subject has four to six images, with three examples per subject used for training and the remaining for validation. For each subject, 3 data points are used for the training dataset, and 1 to 3 data points are used for the validation dataset. To enable the model to distinguish between different subjects, we assign a unique random string to each subject in the prompt. For instance, two different dogs are referenced as “a i0VpE dog” and “a oHwLM dog” in the training and validation sets, respectively.
Training. We use LoRA (DBLP:conf/iclr/HuSWALWWC22:LoRA) to fine-tune stable-diffusion-v1.5 (DBLP:conf/cvpr/RombachBLEO22:StableDiffusion). We apply LoRA to every attention layer in the stable-diffusion-v1.5 model. The basic hyper-parameters settings are listed as follows: minibatch size (4), optimizer (AdamW (DBLP:conf/iclr/LoshchilovH19:AdamW)), initial learning rate ($10^{-4}$), number of steps (10000), and learning rate decay (cosine). For LoRA hyperparameters, we set $r=2$ and $\alpha=2$. The training was performed using the HuggingFace Peft library (peft).
@@ C.1 Comparison with EK-FAC
We compare our proposed VM/FVM approach with EK-FAC (DBLP:journals/corr/abs-2308-03296:EK-FAC) on mislabeled sample detection tasks, following the experimental setup described in Section B.1. The results are presented in Table 7, where * indicates performance based on the best training checkpoints. As shown, our method consistently outperforms the EK-FAC-based influence function in identifying mislabeled data.
@@ C.3 Sharpness-Aware Optimizer
We compare three different sharpness-aware optimizer, including SAM (utilized in our initial approach), ASAM (DBLP:conf/icml/KwonKPC21:ASAM), and F-SAM (DBLP:conf/cvpr/LiZHCH24:F-SAM), on CIFAR-10N Worst and CIFAR-100N Noisy datasets. The results (ROC AUC) are presented in Table 9. As observed, different sharpness-aware optimizers do affect the final results, with F-SAM achieving the best performance. This supports our hypothesis that better flat validation minima can lead to more accurate influence function estimation.
@@ C.4 Inverse Hessian Approximation
To quantitatively discover the impact of the approximation on the inverse Hessian, we conduct experiments on the mislabel detection task. Specifically, we replace the computation of the inverse Hessian $\tilde{H}_{val}^{-1}$ in Equation 21 from diagonal Fisher to LiSSA (DBLP:conf/icml/KohL17:IF) and DataInf (DBLP:conf/iclr/KwonWW024:DataInf).
@@ E.1 The Parameter Change $\tilde{\theta}_{z_{\text{tr}}}-\tilde{\theta}$
Here, for completeness, we derive the parameter change $\tilde{\theta}_{z_{\text{tr}}}-\tilde{\theta}$ in the context of loss minimization. This derivation builds upon the analysis of $\theta^{\star}_{z_{\text{tr}}}-\theta^{\star}$ presented in DBLP:conf/icml/KohL17:IF.
@@ F.1 Influence Function
The use of influence-based diagnostics dates back to seminal works such as cook1982ResidualsInfluenceInRegression. More recently, DBLP:conf/icml/KohL17:IF introduced Influence Functions to large-scale Deep Neural Networks (DNNs), sparking significant research interest. However, subsequent studies have identified challenges in applying IF to DNNs. DBLP:conf/iclr/BasuPF21:IFareFragile highlighted the fragility of IF in deep learning scenarios, while DBLP:conf/nips/BaeNLGG22:IFQA demonstrated that practical influence function estimates often fail to align with leave-one-out retraining in nonlinear networks. Instead, they approximate a different quantity known as the proximal Bregman response function (PBRF).
Still, IF has inspired a wave of research aimed at improving its computational efficiency and scalability for large models. TracIn (DBLP:conf/nips/PruthiLKS20:TraceIn) improves computational efficiency by tracking loss changes using gradient information from multiple training checkpoints. FastIF (DBLP:conf/emnlp/GuoRHBX21:FastIF) accelerates influence estimation by selecting the $k$-nearest training samples around a given validation sample. DBLP:conf/aaai/SchioppaZVS22:ArnoldiIF speed up the inverse Hessian calculation based on Arnoldi iteration (Arnodli). DBLP:journals/corr/abs-2308-03296:EK-FAC proposed Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) to scale Influence Function computations to large language models (LLMs). GEX (DBLP:conf/nips/KimKY23:GEX) alleviates the bilinear constraint of standard Influence Functions and leverages Geometric Ensemble (GE) for more precise influence estimation. DataInf (DBLP:conf/iclr/KwonWW024:DataInf) further enhanced efficiency by deriving a closed-form influence estimation, making it particularly effective for parameter-efficient fine-tuning techniques such as LoRA (DBLP:conf/iclr/HuSWALWWC22:LoRA).
@@ F.2 Sharpness-Aware Minimization
The relationship between the flatness of local minima and generalization has been extensively explored in prior works (DBLP:conf/icml/DinhPBB17:SharpMinimaGeneralize; DBLP:conf/nips/Li0TSG18:VisualLossLand). More recently, numerous studies have aimed to improve model generalization by explicitly seeking flatter minima (DBLP:conf/iclr/ChaudhariCSLBBC17:EnrtropySGD; DBLP:conf/icml/TsuzukuSS20:NFM; DBLP:conf/iclr/ForetKMN21:SAM). Among these, Sharpness-Aware Minimization (SAM) (DBLP:conf/iclr/ForetKMN21:SAM) has attracted considerable attention. SAM introduces a general training framework that formulates optimization as a min-max problem, encouraging model parameters to lie in neighborhoods with uniformly low loss, thereby achieving state-of-the-art generalization across a wide range of tasks. Building on SAM, subsequent works have advanced the approach by refining the geometric definition of the neighborhood (DBLP:conf/icml/KwonKPC21:ASAM; DBLP:conf/icml/Kim0HH22:FisherSAM), proposing improved surrogate loss functions (DBLP:conf/iclr/ZhuangGYCADtD022:GSAM), developing more adaptive adversarial strategies (DBLP:conf/nips/LiG23:VaSSO), and improving training efficiency (DBLP:conf/iclr/JiangY0K23:AE-SAM).
@@ G.1 Remark on the Sign of Influence
In previous studies, the same training sample can yield positive or negative influence depending on the paper, due to differing definitions of loss change. This ambiguity arises particularly between $\ell(z,\theta_{z})-\ell(z,\theta)$ and $\ell(z,\theta)-\ell(z,\theta_{z})$. For example, DBLP:conf/icml/KohL17:IF define the influence as $\ell(z,\theta_{z})-\ell(z,\theta)$, whereas DBLP:conf/nips/PruthiLKS20:TraceIn define it as $\ell(z,\theta)-\ell(z,\theta_{z})$. In this paper, we adopt the previous approach, measuring loss change as $\ell(z,\theta_{z})-\ell(z,\theta)$.